A simple pendulum has a mass of 0.250 kg and a length of 1.00 m. It



is displaced through an angle of 15.0° and then released. What are (a)



the maximum speed, (b) the maximum acceleration, and (c) the



maximum restoring force? What If? Solve this problem by using the



simple harmonic motion model for the motion of the pendulum, and



then solve the problem more precisely by using more general



principles.

1
Expert's answer
2021-12-15T11:20:06-0500

(a) The speed:


"v=\sqrt{2gL\cos\theta}=4.35\text{ m/s}."

(b) The acceleration:


"a=v^2/r=4.35\text{ m/s}^2."

(c) The maximum restoring force:


"F=m\bigg(g+\frac{v^2}g\bigg)=7.2\text{ N}."